# An algebraic model for rational excisive functors

### v1

@@ 1. Introduction
The basic aim of Goodwillie calculus is to understand the input functor $F$ by analysing its excisive and homogenous approximations. Hence, one studies the structure of the categories of excisive and homogeneous functors in general. In the case of $d$-homogeneous functors from spectra to spectra, Goodwillie’s classification of homogeneous functors says that there is an equivalence of ()limit-from($\infty$-)( ∞ - )categories between the category of $d$-homogeneous functors and spectra with an action of $\Sigma_{d}$ (the group of permutations of $d$ elements)
$\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})\simeq{{\mathsf{Sp}}}^{B\Sigma_{d}},$
@@
The classification of categories of excisive functors is more complicated. For functors between the categories of (pointed) spaces and spectra, Arone and Ching [AC15] provided models in terms of coalgebras over a comonad. Glasman [Gla16] gives a simpler classification of the category $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})$ of $d$-excisive endofunctors of spectra in terms of (spectral) Mackey functors on the category ${\mathsf{Epi}}_{\leq d}$ of epimorphisms of finite sets of cardinality at most $d$. This is analogous to the description of equivariant spectra for a finite group $G$ as (spectral) Mackey functors for the orbit category ${\mathsf{Orb}}_{G}$ of Guillou and May [GM24] and Barwick [Bar17].
Glasman’s result instigated an analogy between equivariant homotopy theory and Goodwillie calculus which was utilised by Arone, Barthel, Heard and Sanders [ABHS24] to provide a tensor triangulated classification of excisive functors through a computation of the Balmer spectrum of the category of (compact) $d$-excisive functors. A key ingredient in this work is the Goodwillie-Burnside ring $A(d)$: the abstract Burnside ring in the sense of Yoshida [Yos87] associated to the category ${\mathsf{Epi}}_{\leq d}$. This is a generalisation of the Burnside ring from equivariant algebra and representation theory. Indeed, inputting the orbit category of a finite group $G$ into Yoshida’s machine produces the familiar Burnside ring.
@@
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}\mathsf{Homog}_{i}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}},$
@@
Results of this form are already present in the literature. For instance, Kuhn [Kuh04] and McCarthy [McC01] used the Kuhn-McCarthy pullback square together with the rational Tate vanishing for finite groups to prove that every $d$-excisive functor splits as a product of its homogeneous layers after rationalization. Our proof of A (an outline is given at the end of the introduction) avoids both the use of the Kuhn-McCarthy pullback square and rational Tate vanishing.
We can provide an even simpler description of $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ using Goodwillie’s classification of homogeneous functors (as described above) and results from rational equivariant stable homotopy theory. Greenlees [Gre06] conjectured that the category of rational $G$-spectra (for $G$ any compact Lie group) has an algebraic model: an algebraic category $\mathcal{A}(G)$ and an equivalence ${{\mathsf{Sp}}}^{G}_{\mathbb{Q}}\simeq\mathcal{A}(G)$. This program has had many successes, for example [BGKS17, GS18, Wim19, BGK21, BHI+22b, BHI+22a], see [BK22] for a survey and introduction to the area. A key method in many of these results is using the idempotents of the rationalised Burnside ring $A_{\mathbb{Q}}(G)$. Applying the fact that the category of rational chain complexes with an action of $\Sigma_{d}$, ${\mathsf{Ch}}(\mathbb{Q}[\Sigma_{d}])$, is an algebraic model for ${{\mathsf{Sp}}}^{B\Sigma_{d}}$ gives the following.
@@
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}{\mathsf{Ch}}(\mathbb{Q}[\Sigma_{i}]),$
@@
We also obtain the following seemingly new result, as 6.5, describing rational homogenous functors as a category of modules. We do not expect this result to hold integrally.
@@
$\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq{\mathsf{Mod}}_{{\mathsf{Fun}}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}}(D_{d}(\Sigma^{\infty}\Omega^{\infty})),$
@@
Taking motivation from equivariant homotopy theory, the first step is to understand the rational Goodwillie-Burnside ring. Following Yoshida, we show (2.4) that the rational Goodwillie-Burnside ring splits
@@
We can represent this ring map as a matrix. Computing the idempotents of the rational Goodwillie-Burnside ring then amounts to inverting this matrix (2.6). In particular, for each $1\leq i\leq d$ we obtain an idempotent $e_{i}$ of the Goodwillie-Burnside ring $A_{\mathbb{Q}}(d)$, corresponding to the projection on the $i$-th factor in the product $\prod_{1\leq i\leq d}\mathbb{Q}$.
By work of Arone, Barthel, Heard and Sanders [ABHS24], the Goodwillie-Burnside ring is the endomorphism ring of the unit of the homotopy category of $d$-excisive functors (5.1), and hence these idempotents lift to the category of rational excisive functors, and induce a splitting of the category of rational excisive functors into an $e_{d}$-local piece and its orthogonal complement a $(1-e_{d})$-local piece.
The final step in the proof is to identify the orthogonal pieces in the splitting. To do this, we compute the kernel of a natural map $A_{\mathbb{Q}}(d)\to A_{\mathbb{Q}}(d-1)$ coming from the inclusion of a subcategory ${\mathsf{Epi}}_{\leq d-1}$ into ${\mathsf{Epi}}_{\leq d}$, showing that it is generated by the top idempotent $e_{d}$ (3.6). This map on rational Goodwillie-Burnside rings corresponds to the universal map $P_{d}F\to P_{d-1}F$ (5.2). It follows that for $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ a rational $d$-excisive functor, peeling off the top idempotent splits the Goodwillie tower of $F$ into two pieces,
$e_{d}$-local$(1-e_{d})$-local
[…]
where the orange box is the $e_{d}$-local component and the yellow box is the $(1-e_{d})$-local component. 6.4 uses this to split the category of $d$-excisive functors into $d$-homogeneous functors and $(d-1)$-excisive functors. The result follows by induction.
@@
Our proof of the rational splitting of the category of excisive functors is very different to those currently in the literature. It relies mainly on the formal properties of the category as developed in [ABHS24], primarily the symmetric monoidal structure of the category and the structure of the rational Goodwillie-Burnside ring. The authors believe there should be further results along these lines, where one has some generalisation of (topological) Mackey functors and can prove a splitting result via properties of idempotents in the rational setting. Indeed, recent work of the first two authors with Hill on incomplete Mackey functors [BHK24], uses a similar splitting approach with rational idempotents and again connects to rational Tate vanishing. The authors believe this approach will be valuable in a setting where the analogue of rational Tate vanishing is not accessible.
We believe that C is new and of some independent interest, providing a description of rational homogeneous functors in terms of a category of modules.
The authors also want to add to the efforts of Arone, Barthel, Heard and Sanders [ABHS24] in advertising the results and methods of Yoshida [Yos87]. In this paper we concentrate on the rational case, and we explicitly describe and draw the categories ${\mathsf{Epi}}_{\leq 2}$, ${\mathsf{Epi}}_{\leq 3}$ and ${\mathsf{Epi}}_{\leq 4}$, calculate a complete set of idempotents for associated rational Goodwillie-Burnside rings in these cases and offer a comparison to the case of ${\mathsf{Orb}}_{G}$ for $G$ the cyclic group of order 6.
@@ Acknowledgements
We are grateful to Luca Pol for helpful conversations relating to this work. The second and third author were supported by the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (Dutch Research Council) Vidi grant no VI.Vidi.203.004.
@@ 2. Rational abstract Burnside rings
In this section, we survey Yoshida’s theory of abstract Burnside rings over the rational numbers [Yos87]. It describes a process of assigning a ring to a finite skeletal category, which in the case of the orbit category of a finite group $G$ recovers (2.7) the rational Burnside ring $A(G)$ from rational equivariant algebra [Yos87, Example 3.3.(a)]. We work rationally since our applications are rational, but the theory holds more generally, see e.g., [Yos87, YOT18].
@@
It follows that for each object $c\in\mathcal{C}$, we have that
@@
By [Yos87, §4.2], the epimorphism matrix is conjugate to a lower triangular matrix, the monomorphism matrix is conjugate to an upper triangular one and the automorphism matrix is diagonal, hence the decomposition of $H$ from 2.3 implies that $\det(H)=\prod_{c\in\mathcal{C}}|\mathsf{Aut}(c)|$, which is non-zero since $\mathsf{Aut}(c)\neq\varnothing$ for every $c\in\mathcal{C}$. It follows that the hom-set matrix is rationally invertible and the Burnside homomorphism is an isomorphism.
@@
The strength in the rational Burnside ring is that the isomorphism of 2.4 allows us to identify idempotents through the process of Möbius inversion, i.e., by inverting the hom-set matrix.
@@
We now examine how functors between finite skeletal categories interact with the abstract Burnside rings. For the remainder of this section, we assume all categories satisfy the conditions of 2.4. A functor $f:\mathcal{C}\to\mathcal{D}$ between finite skeletal categories induces a (unique) ring map
@@
[…]
@@
The functor $i^{\ast}:A_{\mathbb{Q}}(\mathcal{D})\to A_{\mathbb{Q}}(\mathcal{C})$ is the identity on the objects of $\mathcal{C}$ and zero on the objects of $\mathcal{D}$ which are not objects of $\mathcal{C}$, hence each element of $\mathcal{D}\setminus\mathcal{C}$ lies in the kernel of $i^{\ast}$.
For the converse, let $A$ be in the kernel of $i^{\ast}$. Then we may express $A$ as a linear sum $A=\sum_{d\in\mathcal{D}}\alpha_{d}\cdot d$, where $\alpha_{d}\in\mathbb{Q}$. It follows that
$0=i^{\ast}(A)=\sum_{d\in\mathcal{D}}\alpha_{d}\cdot i^{\ast}(d)=\sum_{c\in\mathcal{C}}\alpha_{c}\cdot c,$
and hence
$A=\sum_{c\in\mathcal{C}}\alpha_{c}\cdot c+\sum_{d\in\mathcal{D}\setminus\mathcal{C}}\alpha_{d}\cdot d=0+\sum_{d\in\mathcal{D}\setminus\mathcal{C}}\alpha_{d}\cdot d=\sum_{d\in\mathcal{D}\setminus\mathcal{C}}\alpha_{d}\cdot d\in\langle\mathcal{D}\setminus\mathcal{C}\rangle.\qed$
@@ 3. Epiorbital categories of epimorphisms
We will now provide an example of the abstract Burnside ring machinery of Section 2 by choosing the finite skeletal category to be the (skeleton of the) category ${\mathsf{Epi}}_{\leq d}$ of epimorphisms of finite sets of cardinality at most $d$. In the case $d=2$, there is an equivalence between the category of epimorphisms and the obit category ${\mathsf{Orb}}_{C_{2}}$ of the group $C_{2}$ of two elements. In general, however, the category ${\mathsf{Epi}}_{\leq d}$ is not equivalent to the orbit category of a finite group.
@@
This chosen terminology will be made clear in Section 5 where we will identify (5.2) the abstract Burnside ring with the endomorphism ring of the unit of the $\infty$-category of rational $d$-excisive functors.
@@
We are predominantly interested in understanding the idempotents of the rational Goodwillie Burnside ring $A_{\mathbb{Q}}(d)$. Recall from 2.6 that these idempotents in general are given by
@@
Applying 2.6, the idempotents in the rational Goodwillie-Burnside ring $A_{\mathbb{Q}}(3)$ are given by
@@
Our final low dimensional example highlights the passage between the Goodwillie-Burnside ring $A(d-1)$ and the Goodwillie Burnside ring $A(d)$ through the inclusion of ${\mathsf{Epi}}_{\leq d-1}$ as a full subcategory of ${\mathsf{Epi}}_{\leq d}$, which satisfies the conditions of 2.8.
@@
The inclusion of the full subcategory ${\mathsf{Epi}}_{\leq 3}\hookrightarrow{\mathsf{Epi}}_{\leq 4}$ induces a ring map $A_{\mathbb{Q}}(4)\to A_{\mathbb{Q}}(3)$. This inclusion of the full subcategory satisfies the conditions of 2.8, so the kernel is generated by $[4]\in A_{\mathbb{Q}}(4)$. By our computation of the idempotents above, this is equivalent to the kernel being generated by the “top idempotent” $e_{4}$.
@@
where $|{\mathsf{surj}}(i,j)|$ is the cardinality of the set of surjections from a set with $i$ elements to a set with $j$ elements and $s(i,j)$ is the (signed) Stirling number of the first kind. Observe that the number of surjections is given by, $|{\mathsf{surj}}(i,j)|=j!S(i,j)$, where $S(i,j)$ is the Stirling number of the second kind. In the computation of $H^{-1}$ we make use of the fact that the inverse of $H$ exists since none of the diagonal elements $H_{k,k}=|{\mathsf{surj}}(k,k)|=k!$ are zero. Since $H$ is lower triangular, we have that $H^{-1}$ is also lower triangular, and that the diagonal elements of $H^{-1}$ are the reciprocal of the diagonal elements of $H$. The terms away from the leading diagonal may be computed iteratively, simply using the equation $H^{-1}H=I$. In particular, we see that the “top idempotent” corresponding to $[d]\in{\mathsf{Epi}}_{\leq d}$ is given by
@@
and 2.6 along with our description of $H^{-1}$ gives the other idempotents.
@@
The inclusion of the full subcategory $i_{d-1}:{\mathsf{Epi}}_{\leq d-1}\hookrightarrow{\mathsf{Epi}}_{\leq d}$ induces a ring map $(i_{d-1})^{\ast}:A_{\mathbb{Q}}(d)\to A_{\mathbb{Q}}(d-1)$. This inclusion of the full subcategory satisfies the conditions of 2.8, so the kernel is generated by $[d]\in A_{\mathbb{Q}}(d)$. By our computation of the idempotents above, this is equivalent to the kernel being generated by the “top idempotent” $e_{d}$.

### v2

@@ 1. Introduction
The basic aim of Goodwillie calculus is to understand the input functor $F$ by analysing its excisive and homogenous approximations. Hence, one studies the structure of the categories of excisive and homogeneous functors in general. In the case of $d$-homogeneous functors from spectra to spectra, Goodwillie’s classification of homogeneous functors says that there is an equivalence of ()($\infty$-)categories between the category of $d$-homogeneous functors and spectra with an action of $\Sigma_{d}$ (the group of permutations of $d$ elements)
$\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})\simeq{{\mathsf{Sp}}}^{B\Sigma_{d}},$
@@
The classification of categories of excisive functors is more complicated. For functors between the categories of (pointed) spaces and spectra, Arone and Ching [AC15] provided models in terms of coalgebras over a comonad. Glasman [Gla16] gives a simpler classification of the category $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ of $d$-excisive endofunctors of spectra in terms of (spectral) Mackey functors on the category ${\mathsf{Epi}}_{\leq d}$ of epimorphisms of finite sets of cardinality at most $d$. This is analogous to the description of equivariant spectra for a finite group $G$ as (spectral) Mackey functors for the orbit category ${\mathsf{Orb}}_{G}$ of Guillou and May [GM24] and Barwick [Bar17].
Glasman’s result instigated an analogy between equivariant homotopy theory and Goodwillie calculus which was utilised by Arone, Barthel, Heard and Sanders [ABHS25] to provide a tensor triangular classification of excisive functors through a computation of the Balmer spectrum of the category of (compact) $d$-excisive functors. A key ingredient in this work is the Goodwillie-Burnside ring $A(d)$: the abstract Burnside ring in the sense of Yoshida [Yos87] associated to the category ${\mathsf{Epi}}_{\leq d}$. This is a generalisation of the Burnside ring from equivariant algebra and representation theory. Indeed, inputting the orbit category of a finite group $G$ into Yoshida’s machine produces the familiar Burnside ring.
@@
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}\mathsf{Homog}^{i}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}},$
@@
Results of this form are already present in the literature. For instance, Kuhn [Kuh04] and McCarthy [McC01] used the Kuhn-McCarthy pullback square together with the rational Tate vanishing for finite groups to prove that every $d$-excisive functor splits as a product of its homogeneous layers after rationalization. Our proof of ˜A (an outline is given at the end of the introduction) avoids both the use of the Kuhn-McCarthy pullback square and rational Tate vanishing.
We can provide an even simpler description of $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ using Goodwillie’s classification of homogeneous functors (as described above) and results from rational equivariant stable homotopy theory. Greenlees [Gre06] conjectured that the category of rational $G$-spectra (for $G$ any compact Lie group) has an algebraic model: an algebraic category $\mathcal{A}(G)$ and an equivalence ${{\mathsf{Sp}}}^{G}_{\mathbb{Q}}\simeq\mathcal{A}(G)$. This program has had many successes, for example [BGKS17, GS18, Wim19, BGK21, BHI+22b, BHI+22a], see [BK22] for a survey and introduction to the area. A key method in many of these results is using the idempotents of the rationalised Burnside ring $A_{\mathbb{Q}}(G)$. Applying the fact that the category of rational chain complexes with an action of $\Sigma_{d}$, ${\mathsf{Ch}}(\mathbb{Q}[\Sigma_{d}])$, is an algebraic model for ${{\mathsf{Sp}}}^{B\Sigma_{d}}$ gives the following.
@@
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}{\mathsf{Ch}}(\mathbb{Q}[\Sigma_{i}]),$
@@
We also obtain the following seemingly new result, as ˜6.5, describing rational homogenous functors as a category of modules. We do not expect this result to hold integrally.
@@
$\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq{\mathsf{Mod}}_{{\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}}(D_{d}(\Sigma^{\infty}\Omega^{\infty})),$
@@
Taking motivation from equivariant homotopy theory, the first step is to understand the rational Goodwillie-Burnside ring. Following Yoshida, we show (˜2.4) that the rational Goodwillie-Burnside ring splits
@@
We can represent this ring map as a matrix. Computing the idempotents of the rational Goodwillie-Burnside ring then amounts to inverting this matrix (˜2.6). In particular, for each $1\leq i\leq d$ we obtain an idempotent $e_{i}$ of the Goodwillie-Burnside ring $A_{\mathbb{Q}}(d)$, corresponding to the projection on the $i$-th factor in the product $\prod_{1\leq i\leq d}\mathbb{Q}$.
By work of Arone, Barthel, Heard and Sanders [ABHS25], the Goodwillie-Burnside ring is the endomorphism ring of the unit of the homotopy category of $d$-excisive functors (˜5.1), and hence these idempotents lift to the category of rational excisive functors, and induce a splitting of the category of rational excisive functors into an $e_{d}$-local piece and its orthogonal complement a $(1-e_{d})$-local piece.
The final step in the proof is to identify the orthogonal pieces in the splitting. To do this, we compute the kernel of the natural map $A_{\mathbb{Q}}(d)\to A_{\mathbb{Q}}(d-1)$ coming from the inclusion of a subcategory ${\mathsf{Epi}}_{\leq d-1}$ into ${\mathsf{Epi}}_{\leq d}$, showing that it is generated by the top idempotent $e_{d}$ (˜3.6). This map on rational Goodwillie-Burnside rings corresponds to the universal map $P_{d}F\to P_{d-1}F$ (˜5.2). It follows that for $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ a rational $d$-excisive functor, peeling off the top idempotent splits the Goodwillie tower of $F$ into two pieces,
$e_{d}$-$(1-e_{d})-local$
$F{{P_{d-1}F}}{\cdots}$$P_{1}F$$D_{d}F$$D_{d-1}F$$D_{1}F$$\scriptstyle{\simeq}$
where the orange box is the $e_{d}$-local component and the yellow box is the $(1-e_{d})$-local component. ˜6.4 uses this to split the category of $d$-excisive functors into $d$-homogeneous functors and (d1)(d-1)-excisive functors. The result follows by induction.
@@
Our proof of the rational splitting of the category of excisive functors is very different to those currently in the literature. It relies mainly on the formal properties of the category as developed in [ABHS25], primarily the symmetric monoidal structure of the category and the structure of the rational Goodwillie-Burnside ring. The authors believe there should be further results along these lines, where one has some generalisation of (topological) Mackey functors and can prove a splitting result via properties of idempotents in the rational setting. Indeed, recent work of the first two authors with Hill on incomplete Mackey functors [BHK24], uses a similar splitting approach with rational idempotents and again connects to rational Tate vanishing. The authors believe this approach will be valuable in a setting where the analogue of rational Tate vanishing is not accessible.
We believe that ˜C is new and of some independent interest, providing a description of rational homogeneous functors in terms of a category of modules.
The authors also want to add to the efforts of Arone, Barthel, Heard and Sanders [ABHS25] in advertising the results and methods of Yoshida [Yos87]. In this paper we concentrate on the rational case, and we explicitly describe and draw the categories ${\mathsf{Epi}}_{\leq 2}$, ${\mathsf{Epi}}_{\leq 3}$ and ${\mathsf{Epi}}_{\leq 4}$, calculate a complete set of idempotents for associated rational Goodwillie-Burnside rings in these cases and offer a comparison to the case of ${\mathsf{Orb}}_{G}$ for $G$ the cyclic group of order 6.
@@ Acknowledgements
We are grateful to Luca Pol for helpful conversations relating to this work and the anonymous referee for their useful comments. The authors would like to thank the Fondation des Treilles for its support and hospitality during the workshop “Modelling symmetries using algebra”, where work on this paper was undertaken. The second and third author were supported by the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (Dutch Research Council) Vidi grant no VI.Vidi.203.004. In the final stages of preparing this article for publication, the third author was supported by the Engineering and Physical Sciences Research Council under grant number EP/Z534705/1.
@@ 2. Rational abstract Burnside rings
In this section, we survey Yoshida’s theory of abstract Burnside rings over the rational numbers [Yos87]. It describes a process of assigning a ring to a finite skeletal category, which in the case of the orbit category of a finite group $G$ recovers (˜2.7) the rational Burnside ring $A(G)$ from rational equivariant algebra [Yos87, Example 3.3.(a)]. We work rationally since our applications are rational, but the theory holds more generally, see e.g., [Yos87, YOT18].
@@
If $\mathcal{C}$ is a finite skeletal category, it follows that for each object $c\in\mathcal{C}$
@@
By [Yos87, §4.2], the epimorphism matrix is conjugate to a lower triangular matrix, the monomorphism matrix is conjugate to an upper triangular one and the automorphism matrix is diagonal, hence the decomposition of $H$ from ˜2.3 implies that $\det(H)=\prod_{c\in\mathcal{C}}|\mathsf{Aut}(c)|$, which is non-zero since $\mathsf{Aut}(c)\neq\varnothing$ for every $c\in\mathcal{C}$. It follows that the hom-set matrix is rationally invertible and the Burnside homomorphism is an isomorphism.
@@
The strength in the rational Burnside ring is that the isomorphism of ˜2.4 allows us to identify idempotents through the process of Möbius inversion, i.e., by inverting the hom-set matrix.
@@
We now examine how functors between finite skeletal categories interact with the abstract Burnside rings. For the remainder of this section, we assume all categories satisfy the conditions of ˜2.4. A functor $f:\mathcal{C}\to\mathcal{D}$ between finite skeletal categories induces a (unique) ring map
@@
$A_{\mathbb{Q}}(\mathcal{D})$$\mathbb{Q}^{\mathcal{D}}$$A_{\mathbb{Q}}(\mathcal{C})$$\mathbb{Q}^{\mathcal{C}}$$\scriptstyle{\phi_{\mathcal{D}}}$$\scriptstyle{f^{\ast}}$$\scriptstyle{f^{\ast}}$$\scriptstyle{\phi_{\mathcal{C}}}$
@@
The functor $i^{\ast}:A_{\mathbb{Q}}(\mathcal{D})\to A_{\mathbb{Q}}(\mathcal{C})$ is the identity on the objects of $\mathcal{C}$ and zero on the objects of $\mathcal{D}$ which are not objects of $\mathcal{C}$, hence each element of $\mathcal{D}\setminus\mathcal{C}$ lies in the kernel of $i^{\ast}$. The proof is then complete by observing that the dimension of the kernel and the dimension of the free $\mathbb{Q}$-module $\langle\mathcal{D}\setminus\mathcal{C}\rangle$ coincide. ∎
@@ 3. Epiorbital categories of epimorphisms
We will now provide an example of the abstract Burnside ring machinery of Section˜2 by choosing the finite skeletal category to be the (skeleton of the) category ${\mathsf{Epi}}_{\leq d}$ of epimorphisms of finite sets of cardinality at most $d$. In the case $d=2$, there is an equivalence between the category of epimorphisms and the obit category ${\mathsf{Orb}}_{C_{2}}$ of the group $C_{2}$ of two elements. In general, however, the category ${\mathsf{Epi}}_{\leq d}$ is not equivalent to the orbit category of a finite group.
@@
This terminology follows [ABHS25, Section 8]. In Section˜5 we will identify (˜5.2) the abstract Burnside ring with the endomorphism ring of the unit of the $\infty$-category of rational $d$-excisive functors.
@@
We are predominantly interested in understanding the idempotents of the rational Goodwillie Burnside ring $A_{\mathbb{Q}}(d)$. Recall from ˜2.6 that these idempotents in general are given by
@@
Applying ˜2.6, the idempotents in the rational Goodwillie-Burnside ring $A_{\mathbb{Q}}(3)$ are given by
@@
Our final low dimensional example highlights the passage between the Goodwillie-Burnside ring A(d1)A(d-1) and the Goodwillie Burnside ring $A(d)$ through the inclusion of ${\mathsf{Epi}}_{\leq d-1}$ as a full subcategory of ${\mathsf{Epi}}_{\leq d}$, which satisfies the conditions of ˜2.8.
@@
The inclusion of the full subcategory ${\mathsf{Epi}}_{\leq 3}\hookrightarrow{\mathsf{Epi}}_{\leq 4}$ induces a ring map $A_{\mathbb{Q}}(4)\to A_{\mathbb{Q}}(3)$. This inclusion of the full subcategory satisfies the conditions of ˜2.8, so the kernel is generated by $4]\in A_{\mathbb{Q}}(4)$. By our computation of the idempotents above, this is equivalent to the kernel being generated by the “top idempotent” $e_{4}$.
@@
where $|{\mathsf{surj}}(i,j)|$ is the cardinality of the set of surjections from a set with $i$ elements to a set with $j$ elements and $s(i,j)$ is the (signed) Stirling number of the first kind. Observe that the number of surjections is given by, $|{\mathsf{surj}}(i,j)|=j!S(i,j)$, where $S(i,j)$ is the Stirling number of the second kind. In the computation of $H^{-1}$ we make use of the fact that the inverse of $H$ exists since none of the diagonal elements $H_{k,k}=|{\mathsf{surj}}(k,k)|=k!$ are zero. Since $H$ is lower triangular, we have that $H^{-1}$ is also lower triangular, and that the diagonal elements of $H^{-1}$ are the reciprocal of the diagonal elements of $H$. The terms away from the leading diagonal may be computed iteratively, simply using the equation $H^{-1}H=I$. In particular, we see that the “top idempotent” corresponding to $d]\in{\mathsf{Epi}}_{\leq d}$ is given by
@@
and ˜2.6 along with our description of $H^{-1}$ gives the other idempotents.
@@
The inclusion of the full subcategory $i_{d-1}:{\mathsf{Epi}}_{\leq d-1}\hookrightarrow{\mathsf{Epi}}_{\leq d}$ induces a ring map $(i_{d-1})^{\ast}:A_{\mathbb{Q}}(d)\to A_{\mathbb{Q}}(d-1)$. This inclusion of the full subcategory satisfies the conditions of ˜2.8, so the kernel is generated by $d]\in A_{\mathbb{Q}}(d)$. By our computation of the idempotents above, this is equivalent to the kernel being generated by the “top idempotent” $e_{d}$.